During rush hour, Fernando can drive 40 miles using the side roads in the same time that it takes to travel 30 miles on the freeway. If Fernando's rate on the side roads is 7 mi/h faster than his rate on the freeway, find his rate on the side roads.

Answer by scott8148(6628)

let x=rate on side roads ... in the same time, he can only go 3/4 as far on the freeway, so the freeway rate is (3/4)x

(3/4)x+7=x ... 7=(1/4)x ... 28=x

using t=d/r ... the times are the same so 40/r=30/(r-7)

cross multiplying gives 40r-280=30r ... 10r=280 ... r=28